Церковний рiк Страстi Христовi

30. Агнець Божий, чистий i Святий

Nikolaus Decius, 1485-1546 O LAMM GOTTES, UNSCHULDIG Пер. Т. Коковського, 1956 Nikolaus Decius, 1485-1546

[figure: 4 4; 4 4; {1. Бо \atop 2. Аг}; {- жий \atop -нець,}; 2; {Син, \atop що}; {чис \atop був}; {тий \atop роз}; 3; {й \atop - п’я}; свя; -; {- \atop -}; {- \atop -}; 4; {тий, \atop тий,}; {пi \atop Чий}; 5; {- шов \atop пос}; {на \atop - лух}; 6; хрест нам; {у \atop на}; 7; {му \atop - у}; -; {- \atop -}; {ках. \atop ка.}]

[figure: 8; По; 9; - ми; - луй; 10; же; нас,; 11; о; I; 12; - су; -; -; -; 13; се!; 14; О; I; -]

[figure: 15; су; -; -; -; 16; се!; 17; Вiн; 18; всi; грi; 19; - хи; по; 20; - нiс; за]

[figure: 21; нас,; та; 22; вiд; грi; хiв; всiх; 23; нас; спас.; 24; Свiй; 25; мир; зi; 26; - шли; нам,; 27; о; I; -]

[figure: 28; су; -; -; -; 29; се!; 30; О; I; 31; - су; -; -; -; 32; се!; 33; А; - мiнь.]

†Павло Богмат MusiX TEX